# The building blocks of software work explain coding careers and language popularity

### v1

@@ 3 Discussion
The growing complexity of the labor market has made measuring worker capabilities both more difficult and more important (Anderson, 2017). Here we demonstrated that data on individual problem solving can be used to generate a useful fine-grained taxonomy of tasks people do in software-related work, a key industry in the digital economy. These tasks or building blocks of software work can be given coherent economic values and capture co-occurrence relationships reflected in real world job advertisements. We use the tasks to observe individual learning, both within tasks and entrance into new ones. Moreover, the fine-grained detail of the task space allows us to explain an important macro-system outcome: the rise of Python as the leading all-purpose programming language. Our explanation for Python’s success is that it enables people to enter new tasks which have higher economic value. Ultimately, our results highlight how deepening our understanding of the microstructure of tasks offers new ways to interpret macro-level phenomena such as technology diffusion, wage dispersion, and the evolution of industry-wide practices.
A core insight from our results is that tasks exist in a space of related capabilities, echoing network-based accounts of human capital (Anderson, 2017). On the one hand, tasks are well-defined as distinctive categories of specific kinds of work, as evidenced by their distinct labor market values and that effective learning happens by practice within tasks (Arrow, 1962). On the other hand, tasks also have meaningful relationships with each other, evidenced by their non-random co-occurrence within job ads and observed patterns of individuals learning new tasks. The resulting networked taxonomy of software tasks demonstrates its value by providing a micro-to-macro explanation for the rise of the Python programming language: Python enables users to enter more valuable tasks.
@@
Our findings connect with the literature on skill complexity, labor market transitions, and the organization of work, complementing studies that have emphasized how granular task analysis can reveal hidden dynamics in career mobility, wage disparities, and skill complementarities (Autor et al., 2003; Anderson, 2017; Alabdulkareem et al., 2018; del Rio-Chanona et al., 2021). Previous research however, tends to use standardized task classifications such as O*NET to define tasks and applies them to curated job descriptions, limiting the granularity and responsiveness of task taxonomies. In contrast, by extracting tasks directly from real-world problem-solving interactions on Stack Overflow, our approach captures the immediate and practical realities of skill application at a global scale. In rapidly changing sectors such as software engineering, tasks derived from activities can explain surprising developments like the rise of the Python programming language as a technology which enables diversification into more valuable tasks. This complements existing work on task-relatedness and collective learning.
@@
We highlight three possible applications of our method and taxonomy. First we have demonstrated that microdata on activity can be fruitfully aggregated into tasks to describe skill acquisition, development and transitions. Tasks from any sector in which large-scale fine-grained activity on individual behavior at work is available could be categorized in a similar manner. Second, employers in such sectors can use our approach to identify skills related to tasks their employees actually carry out, enabling better labor market matching outcomes or internal skill development programs. Finally, educators can better understand the labor market landscape of the software industry using our taxonomy.
@@
There are several limitations to our study, which may be tackled in future work. First, we have focused exclusively on programming-related tasks. However, software developers need to be able to carry out many other tasks, such as managing teams or reporting results. Although SO contains some postings about organizational aspects of software development, its emphasis is on technical subjects. To explore how programming tasks are combined with other types of tasks in software jobs, future studies could use job ads to compare the co-occurrence of our programming tasks with tasks related to different kinds of soft skills.
@@
Second, although the question-answer structure makes SO ideal for identifying software tasks, it may be less suited for descriptions of macro-level trends and individual career paths. In Appendix E, we compare user numbers on SO to GitHub. This analysis suggests that in terms of capturing software developers, both datasets are closely aligned. However, it is impossible to judge the representativity of SO of overall software activity at the level of tasks, especially as Stack Overflow is thought to be declining (del Rio-Chanona et al., 2024). In Appendix E, we therefore rescale the activity in different programming languages such that they reflect their presence on GitHub. However, in future work, we plan to define software tasks directly in GitHub.
@@
Third, we opted to create a task taxonomy that is time-invariant. However, an analysis of how job tasks change in software development would provide a more vivid picture of the software industry. Such an analysis could also expand the size of our job ad dataset. Although sufficient for the validation exercise we conducted, a larger dataset would allow us to see how tasks in software development and their value change over time.
@@
Notwithstanding these limitations, our study demonstrates how question-and-answer data can inform us about the evolution of work by quantifying the abstract notion of a task. The application to the software sector sheds light on novel aspects of high skill labor markets at high granularity, at scale, and in near real-time. Moreover, programming tasks are rapidly diffusing beyond pure software development firms and understanding these tasks and their evolution is pivotal to understanding the future of work in an increasingly digital economy.
4 Methods
4.1 Data
Stack Overflow Posts:
Our main dataset is the corpus of questions and answers (together “posts”) on Stack Overflow from the site’s launch in August 2008 to June 2023. The 23 million questions in this dataset are labeled by tags from a curated system of about 66,000 tags referring to software concepts and technologies such as key-value, scipy and svm. For each tag, SO provides a textual description, typically consisting of a few sentences. Tags are heavily moderated by humans and automated processes, and are deduplicated through a dictionary of synonyms. We focus on tags that are used at least 1,000 times. We split tags into a two groups: 5,083 general tags and 282 tags that refer to programming languages, defined broadly. For the latter, we merge different versions of the same programming language, converting, for instance python-3.x into the generic tag python. When we define tasks, we use all available posts, with answers inheriting the tags of the questions they answer. In all other analyses, we only consider answers from users with a total of at least 10 answer posts.
Stack Overflow Develop Survey:
Stack Overflow provides access to yearly surveys of its users, the Stack Overflow Developer Survey, with anonymous information on salaries, as well as on tools and languages used that can often be linked to tags in SO. We use the survey data from year 2023, which contains around 12,000 participants who report their salaries to estimate task values.
HackerNews Job Ads:
We scraped job advertisements from “hnhiring.com”, an index of jobs from the website Hacker News’ “Who is Hiring?” posts. This yielded a total of 50,998 job advertisements. We use approximately 46,000 ads to study task co-occurrence and 3,949 jobs to test the ability of tasks to predict salaries.
4.2 Building software tasks
4.2.1 Identifying programming languages
To identify tags that refer to programming languages, we rely on wiki pages, which are available for most common tags. We first filter the pages to find tags that include “language” or “framework” among their keywords. We then manually review all the tags to select those that refer to programming languages, yielding a total of 282 programming languages.
Some programming tags refer to specific versions of a language, such as “python-3.x” or “python-3.7”. When calculating the nestedness of the task-language matrix, we merge all tags referring to different versions of the same programming language. Yet, there is some ambiguity to what counts as a programming language, such as frameworks and developer environments. We remove such tags when calculating nestedness, basing this estimate on a task-language matrix with 114 programming languages.
4.2.2 Identifying software tasks
To identify canonical software tasks, we express the relation between questions and their tags as a bipartite network with 18,154,593 question nodes that connect to 5,083 tags through 33,378,590 edges. We then apply a bipartite Stochastic Block Model (SBM) to identify communities of tags.
SBMs essentially estimate a coarsened version of a network. To do so, they identify communities of nodes and parameters that represent parsimonious descriptions of edge distributions within and between communities. To sharpen the boundaries of inferred communities, we remove $20\%$ of tags that have the weakest own-community attachment (see Appendix section A). Dropping communities with fewer than 3 tags, we retain 237 tasks, connected to 4,054 tags, out of originally 247 detected communities.
4.3 User task vectors
We use the tasks extracted from the question-tag network to describe the activity of individual SO users as a 237-dimensional vector. Entries of this vector count how many answers the user has provided to questions that are labeled with tags that are associated with each corresponding task. Specifically, we express the user’s task experience in time period $t$ as:
$\vec{T}_{u}^{t}=\begin{bmatrix}X_{u,1}^{t}\\ ...\\ X_{u,\theta}^{t}\\ ...\\ X_{u,N_{\theta}}^{t}\\ \end{bmatrix},$ (1)
where $N_{\theta}=237$ the total number of tasks and $X_{u,\theta}^{t}$ is the number of answers posted in period $t$ by user $u$ to questions labeled with a tag belonging to the community of tags associated with task $\theta$.
4.4 Constructing a Task Space
To calculate the relatedness between tasks, we analyze which tasks are often associated with the same users. To do so, we collect the task vectors of all users in an $(N_{\theta}\times N_{u})$ matrix T, with $N_{u}$ the number of users in the dataset with at least 10 answers and $N_{\theta}=237$ the total number of tasks. Next, we ask how often two tasks co-occur within the same users by multiplying $(\textbf{T}>0)$ with its transpose: $(\textbf{T}>0)(\textbf{T}>0)^{\prime}$. Finally, we calculate the Pointwise Mutual Information ($\mathrm{PMI}$) of probability $p_{\theta,\kappa}$ that task $\theta$ and $\kappa$ co-occur in the same user to capture the amount of (information-theoretic) surprise involved in observing the joint probability $p_{\theta,\kappa}$ against a benchmark where draws of tasks $\theta$ and $\kappa$ are independent:
$\mathrm{PMI}_{\theta,\kappa}=\log\frac{p_{\theta,\kappa}}{p_{\theta}p_{\kappa}},$ (2)
where $p_{\theta}$ and $p_{\kappa}$ are marginal probabilities, i.e., $p_{\theta}=\sum_{\gamma}p_{\theta,\gamma}$ and $p_{\kappa}=\sum_{\mu}p_{\mu,\kappa}$.
To estimate $\mathrm{PMI}$ values and their credible intervals, we apply a Bayesian statistical framework developed by (van Dam et al., 2023).
The task space is based on a weighted network where nodes represent tasks and edges the relatedness between them. Because our focus is on relatedness as expressed in a surprisingly large observed probability that two tasks co-occur, we ignore tasks that co-occur less frequently than our random null model predicts by setting all negative $\mathrm{PMI}$ values to 0. To identify clusters in the task space, we embed the relatedness network in a 5-dimensional space using UMAP (McInnes et al., 2018). Next, we cluster sets of contiguous tasks using HDBSCAN (McInnes et al., 2017). These clusters are represented in colors of Fig. 1b. The nodes’ coordinates in this figure are determined by a UMAP embedding of the relatedness network in a 2-dimensional space.
4.5 Task Values
We estimate the value of a task based on salary information from the Stack Overflow developers survey. We focus on the 12,000 US participants that report a salary in 2023. Respondents list the technologies they use, which often correspond to tags in the SO platform. For each SO user, we use the top 300 most similar survey respondents in terms of overlap in tags and technologies to estimate user salary. We use the weighted average salary reported by the most similar respondents, weighting by tag similarity to the SO user. To assign values to tasks, we then calculate the weighted average salary of the users that carry out that task, using the prevalence of the task in users’ task vectors for the years 2018-2023 (the period in which we collect job ads) as weights:
$V_{\theta}$ $=\sum_{u}\frac{X_{u,\theta}^{2018,2023}}{\sum_{u^{\prime}}X_{u^{\prime},\theta}^{2018,2023}}V_{u}$ (3)
where $M_{u,\rho}$ is the number of overlapping tags between SO user $u$ and one of the 300 closest survey respondents, $\rho\in R_{u}$, and $V_{\rho}$ is the salary of survey respondent $\rho$, such that $V_{u}$, the match-weighted average salary of survey respondents matched to user $u$, can be interpreted as a proxy for the user’s salary.
4.6 Job Advertisements
The data on job ads come from the website HackerNews, and were scraped from the aggregator site hnhiring.com. They represent all 50,998 job ads posted to the website between January 2018 and May 2024. These ads are unstructured job descriptions sometimes including salary and expected work activities (tasks). We prompted an LLM (ChatGPT-3.5) to extract salary information and task requirements, which yields about 46,000 jobs listing at least 3 tasks, and 3,949 jobs that contain at least 1 task and salary information. When ads describe more than one position, the LLM provides information on each job separately.
We first relate the task requirements of job ads to the 237 SO tasks. We embed both job ad tasks and SO tags in a 384-dimensional vector space using sBERT (Reimers and Gurevych, 2019). To describe an SO task in this vector space, we calculate the embedding of its label. To increase accuracy, we use somewhat longer labels of up to 30 words generated by ChatGPT-4.0. Finally, we match each task requirement of a job ad to the closest SO tasks based on cosine similarity between the embeddings of the job-ad task and of the SO task. When we cannot find any SO task with a cosine similarity of at least 0.3, we drop the task requirements. This yields a 237-dimensional vector of zeros and ones that describe each job’s requirements in terms of our SO tasks. In Appendix C.3, we demonstrate robustness of our findings to alternative methods and thresholds.
We use these data in two ways to test the validity of our SO tasks. First, we ask whether we can predict the task composition of a job. To do so, we take the SO-task vectors associated with each job ad and stack these vectors into a dataset that contains for each combination of an SO task and an advertised job whether or not the task is required by the job. Next, we mask 40% of observations in this dataset and calculate the density of required tasks in the job around the masked tasks:
$\mathrm{D}_{\theta,j}=\bar{\mathrm{R}}_{\theta,j}\mathrm{T}_{j},$ (5)
where $\bar{}$ denotes row-normalization and $\mathrm{T_{j}}$ and $\mathrm{R}_{\theta,j}$ are the unmasked portions of job $j$’s task vector and the corresponding rows of relatedness matrix $\mathrm{R}$, namely $\mathrm{R}_{\theta,i}=\frac{\mathrm{R}_{\theta,i}}{\sum_{\tau}\mathrm{R}_{\theta,\tau}}$. $\mathrm{D}_{\theta,j}$ denotes the average relatedness of (masked) task $\theta$ to all unmasked tasks in job $j$.
We plot how the estimated probability that a job requires a particular (masked) task changes with that task’s density in the job’s unmasked task requirements. To do so, we bin all masked observations in 10 equally sized bins and calculate relative frequencies with their 95% confidence intervals. These estimated probabilities are plotted on the vertical axis of Fig. 2b, against the bin’s average task density on the horizontal axis.
Our second validation focuses on the set of advertised jobs that also post salary information. It estimates the association between the average value of a job’s SO tasks and the posted wage in the job ad. We report on this in SI, section C.2.
4.7 Analysis of SO users
To study user behavior, we randomly divide the population of users into two equally sized samples. We use the first sample, $S_{1}$, to construct a task relatedness matrix from co-occurrence patterns of tasks in users. The second sample, $S_{2}$, is used to estimate how relatedness affects user activity on SO. To do so, we collect information from all answer posts by these users within rolling two-year intervals. That is, we construct experience vectors $\vec{T}_{u}^{t}$, where $t$ denotes an interval that stretches across the two calendar years preceding year t. For instance, $\vec{T}_{u}^{2018}$ refers to the answers user $u$ provides between January 1, 2016 and December 31, 2017.
4.7.1 Task diversification
Analogously to eq. (5), we use sample $S_{2}$ to calculate for a given user, $u$, and year $t$, the experience density around each task:
$\mathrm{D}^{t}_{u}=\bar{\mathrm{R}}_{S_{1}}\mathrm{T}^{t}_{u},$ (6)
where $\bar{\mathrm{R}}_{S_{1}}$ is the row-normalized relatedness matrix, constructed using users in sample $S_{1}$.
Next, we ask which new tasks users are most likely to enter. To do so, we take all tasks in which user $u$ had no activity in the two calendar years preceding $t$. These tasks represent the set of tasks that user $u$ can enter. We tag all such tasks with a 1 whenever the user posts at least one answer related the the task in year $t$, and with 0 otherwise. Next, we divide all user-task combinations at risk of entry into 10 equally sized bins and estimate the probability that users do enter these tasks in year $t$. Figure 3b shows how these estimated probabilities rise with the user’s experience density around a task. 95% confidence intervals are approximated with a normal distribution, using the following estimate of the standard error of the mean:
$\hat{\sigma}_{b}=\frac{1}{\sqrt{n_{b}}}\sqrt{\hat{\pi}_{b}(1-\hat{\pi}_{b})},$ (7)
where $n_{b}$ the number of observations in bin $b$ and $\hat{\pi}=\frac{d_{b}}{n_{b}}$, where $d_{b}$ the number of entered tasks in bin $b$, the relative frequency with which tasks in bin $b$ are entered.
4.7.2 Voting
To study of a user’s proficiency and learning in a task, we analyze the votes the user’s answers receive. Votes are social feedback from other users indicating appreciation for a post. However, the sample of answers that users post is highly selected. First, before deciding to post an answer, users will look at existing answers to a question to see if they can improve or add to them. To avoid sample selection bias from this, we limit ourselves to the first recorded answers for each question. Second, users will post answers to question for which they perceive themselves to be sufficiently knowledgeable. We address this type of sample selection bias in Appendix D.2, where we use variation in the user base’s activity patterns across the 24-hour daily cycle.
To determine the importance of users’ task experience, we estimate its effect on the upvotes their answers receives. To do so, we estimate the following regression model, with outcomes reported in Fig. 3c:
where $X_{\theta,u(a)}$ is user $u(a)$’s experience in terms of the number of answers provided to questions involving task $\theta$ in the two calendar years preceding the calendar year in which answer $a$ was provided, $A_{q(a)}$ the number of answers provided to question $q(a)$, $Q_{a}$ the set of answers provided to question $q(a)$, $v_{i}$ the number of upvotes that answer $i$ received and $\eta_{a}$ a disturbance term. Furthermore, the dependent variable $y_{a,\theta}$ is either a binary variable that indicates whether answer $a$ is a question’s top answer or $\log(v_{a}+1)$, the logarithm of the number of upvotes answer $a$ receives, augmented by 1 to avoid $\log(0)$ issues. Note that if an answer is associated with multiple tasks, we replicate the observation accordingly.
4.7.3 Task adoption analysis
To study which new tasks users adopt, we estimate the effect of task value and density on new task entries, estimating the following regression model:
$y_{u,\theta}=\beta_{V}\log(V_{\theta})+\beta_{D}D_{u,\theta}+\sigma_{u,\theta},$ (9)
where $y_{u,\theta}$ is a binary variable that encodes whether or not user $u$ adopted task $\theta$, $\log(V_{\theta})$ is the logarithm of the task’s value as defined in eq.( 3), and $D_{u,\theta}$ the user $u$’s density around task $\theta$ as defined in eq. (6) before. We estimate this model once with data from all programming languages and once with data on answers related to Python only. Findings are summarized in Fig. 3d, with detailed results described in Appendix D.3.
Acknowledgements
XF, JW, SD and FN acknowledge financial support from the Austrian Research Agency (FFG), project #873927 (ESSENCSE). JW acknowledges support from the Hungarian National Scientific Fund (OTKA FK 145960). SD acknowledges financial support from the Dutch Research Council (NWO) under the REWIRE project, grant number KICH1.MV02.22.010.
@@ Appendix A Stochastic block models
SBMs represent generative models for networks, that coarsen networks by modeling groupings of nodes and the edge distributions between and within these groupings. These models typically postulate a stochastic data generating process of the network that produces communities and the edges among nodes within and between these communities. Under this assumption, the observed network can be described as sampled from this process.
More precisely, given an $N$-nodes network, $G$, and partition or community structure ${\bf c}$ assigning nodes to different groups, the probability of generating the graph $G$ from its partition ${\bf c}$ is:
@@
where $\theta$ is a set of parameters that govern how nodes connect to one another, depending on the communities they belong to. The inference procedure inverses this relation, such that it recovers the probability that a specific generative model is correct, given the observed network $G$. Thus, the probability of observing the configuration given observed network $G$ with model parameters that observe prior probabilities $P(\theta,{\bf c})$ is:
@@
Under the assumption that the generative model has only one set of particular parameters for network $G$, the equation simplifies to:
@@
In the bipartite implementation of this model, priors are chosen such that nodes of the same layer cannot link to one another Gerlach et al. [2018]. Applying this bipartite SBM to our question-tag network yields communities for both questions and tags. In this research, we are only interested in communities of tags (as opposed to questions) of which we discover 247, containing a total of 5,083 tags.
The procedure yields a community label for each tag, even when there is some ambiguity over which community fits the tag best. For instance, some tags are highly ubiquitous (e.g., pandas in Python related questions) and therefore connect to nodes in various communities. This leads to noise in the community composition. To reduce this noise, we try to remove tags that are only weakly connected to the other tags in their community. In principle, SBM models can identify such nodes using MCMC simulations. However, the size of our question-tag renders this approach computationally infeasible. Therefore, we here use a simpler approach.
In particular, we construct a weighted unipartite tag network, where edge weights denote the intensity with which tags co-occur on the same questions. Next, we label nodes in this tag network according to the community structure we discovered above. For each tag, we now calculate the overrepresentation, $O_{tc}$ of (weighted) links to nodes in each of the 247 communities:
@@
$O_{tc}$ measures how important tag $t$ is to community $c$ compared to all other communities. High values of $O_{ct}$ indicate that the tag is a core part of the community, while low $O_{ct}$ values signal that the tag may be misplaced or only weakly related to community $c$. To focus on core tags, for each community, we remove the $20\%$ of tags with the lowest $O_{ct}$ values. Finally, we also remove communities with fewer than 3 tags, for which the tag composition provides too little information on the associated task. In this way, we arrive at a task space with 237 tasks, composed of a total of 4,054 tags.
Appendix B Task space
In this section, we provide a detailed, annotated version of the task space depicted in Fig. 1b. This annotated task space is shown in Fig. B.1.
Each node is labeled with an example of a programming task provided by ChatGPT-4.0. To do so, we prompt the LLM as follows:
Figure B.1: Fully annotated task space.

### v2

@@ 5 Discussion
The growing complexity of the labor market has made measuring worker capabilities both more difficult and more important. Here we demonstrated that data on individual problem solving can be used to generate a useful, fine-grained taxonomy of skills used in software-related work, a key industry in the digital economy. These skills can be given coherent economic values that help predict wages in job advertisements and be used to guess which questions users provide the best answers to according to feedback from the community. This illustrates the validity and usefulness of the skill taxonomy inside and outside the context of SO.
The fine-grained detail of the skill taxonomy allows us to assess the mapping of programming languages to skills. This reveals a nested landscape of skills, where ubiquitous skills are undemanding in terms of which programming language they require, but specialized skills require not custom-tailored languages but are rather implemented in generalist, highly versatile, languages.
A core insight from our results is that skills exist in a space of related capabilities, echoing network-based accounts of human capital (Anderson, 2017; Alabdulkareem et al., 2018; Stephany and Teutloff, 2024). Constructing a network based on SO postings that connects skills that are often possessed by the same users shows that real-world software development jobs are coherent, requiring skills that are closely related to one another. Moreover, the topology of this network steers how programmers learn: over time SO users acquire new skills that are closely related to the ones they already possess. Another important predictor of skill acquisition is a skill’s value. Somewhat surprisingly, in general, SO users branch into new skills that are of relatively low value. Presumably, these skills are easier to learn, which may explain their modest returns in the first place. However, this behavior is different for Python users, who preferentially acquire high-value skills. Our networked taxonomy of software skills thereby provided a micro-to-macro explanation for the rise of the Python programming language: Python enables users to learn more valuable skills.
This interpretation is consistent with broader indicators of Python’s role in the ecosystem. Python ranks among the most widely used languages in major indices (TIOBE Software BV, 2026), and developer surveys emphasize that it is used across many application areas (JetBrains and Python Software Foundation, 2024). In addition, many AI-assisted data analysis tools solve tasks by executing Python code in a sandboxed environment (OpenAI, ), which makes Python a natural interface between developers and expanding AI-related workflows. Empirical studies have shown that Python is one of the most accessible and ubiquitous languages (Juhász et al., 2026). Taken together, these patterns fit an account in which Python lowers the cost of moving into high-value skill clusters.
@@
Our findings connect with the literature on skill complexity, labor market transitions, and the organization of work, complementing studies that have emphasized how granular human capital analysis can reveal hidden dynamics in career mobility, wage disparities, and skill complementarities (Autor et al., 2003; Anderson, 2017; Alabdulkareem et al., 2018; del Rio-Chanona et al., 2021; Stephany and Teutloff, 2024). Previous research however, tends to use standardized classifications such as O*NET to define skills and applies them to curated job descriptions, limiting the granularity and responsiveness of skill taxonomies. In contrast, by extracting skills directly from real-world problem-solving interactions on SO, our approach captures the immediate and practical realities of skill application at a global scale. The value of this approach is demonstrated in its ability to study and explain phenomena from the micro level (the coherence of job requirements and fine-grained individual learning trajectories) to the macro level (the mapping of skills to programming language and the rapid rise to dominance of Python) linking individual to collective learning dynamics.
@@
We highlight three possible applications of our method and taxonomy. First we have demonstrated that microdata on activity can be fruitfully aggregated into skills to describe skill acquisition, development and transitions. Skills from any sector in which large-scale fine-grained information on individual behavior at work is available could be categorized in a similar manner. Second, in the software sector, employers can use our approach to identify skills related to work activities their employees actually need to perform, enabling better labor market matching outcomes and internal skill development programs. Finally, educators can better understand the labor market landscape of the software industry using our taxonomy.
@@
There are several limitations to our study, which may be tackled in future work. First, we have focused exclusively on programming-related skills. However, software developers also need many other skills, such as management and reporting capabilities. Although SO contains some postings about organizational aspects of software development, its emphasis is on technical subjects. To explore how programming skills are combined with other types of expertise in software jobs, future studies could use job ads to compare the co-occurrence of our programming skills with different kinds of soft skills.
@@
Second, although the question-answer structure makes SO ideal for identifying software work activities and their prerequisite skills, it may be less suited for descriptions of macro-level trends and individual career paths. In Appendix E, we compare user numbers on SO to GitHub. This analysis suggests that in terms of capturing software developers, both datasets are closely aligned. In this Appendix, we therefore also rescale the activity in different programming languages such that they reflect their presence on GitHub. However, it is impossible to judge the representativity of SO at the level of individual programming skills. Moreover, recently, SO usage has strongly declined due to introduction of generative AI and advanced coding assistants (del Rio-Chanona et al., 2024). In future work, we therefore plan to identify software skills directly in code posted on GitHub.
@@
Third, we opted to create a skill taxonomy that is time-invariant. However, an analysis of how programming skills change would provide a more vivid picture of the software industry. Such an analysis could also expand the size of our job ad dataset. Although sufficient for the validation exercise we conducted, a larger dataset would allow us to see how skills in software development and their value change over time.
@@
Notwithstanding these limitations, our study demonstrates how question-and-answer data can inform us about the evolution of software development work by quantifying the abstract notion of a skill. Programming skills are rapidly diffusing beyond pure software development firms and understanding these skills and their evolution is pivotal to understanding the future of work in an increasingly digital economy. Our application contributes to such understanding by shedding light on novel aspects of this at high granularity, at scale, and in near real-time.
Data Availability
Declaration of generative AI use
The authors acknowledge the use of the following AI tools: (a) ChatGPT-3.5 to extract job requirements and salary information from job ads. (b) ChatGPT-4.0 to provide descriptive labels for the skills in our skill taxonomy. These skill labels were used to allow for easy reference to skill examples in tables and in the main text and to link the job requirements extracted from job ads as detailed in point (a) above to our skill taxonomy. In the latter case, findings were corroborated using links that do not rely on the AI-generated skill labels. All prompts are provided in the main content.
@@ Appendix A Stochastic block models
SBMs represent generative models for networks, that coarsen networks by modeling groupings of nodes and the edge distributions between and within these groupings. These models typically postulate a stochastic data generating process of the network that produces communities and the edges among nodes within and between these communities. Observed networks are assumed to be sampled from this process.
More precisely, given an $N$-node network, $G$, and partition or community structure, $\bf c$} that assigns nodes to different groups, the probability of generating graph $G$ from its partition $\bf c$} is:
@@
where $\theta$ is a set of parameters that govern how nodes connect to one another, depending on the communities they belong to. Inference inverses this relation, such that it recovers the probability that a specific generative model is correct, given the observed network $G$. Thus, the probability of observing the configuration, given the observed network $G$ with model parameters that observe prior probabilities $P(\theta,{\bf c})$ is:
@@
Under the assumption that the generative model has only one set of parameters for network $G$, the equation simplifies to:
@@
In the bipartite implementation of this model, priors are chosen such that nodes of the same layer cannot link to one another Gerlach et al. [2018]. Applying this bipartite SBM to our question-tag network yields communities for both questions and tags. In the current paper, we are only interested in communities of tags (as opposed to questions) of which we discover 247, containing a total of 5,083 tags.
The procedure yields a community label for each tag, even when there is some ambiguity over which community fits the tag best. For instance, some tags are highly ubiquitous (e.g., pandas in Python related questions) and therefore connect to nodes in various communities. This leads to noise in the community composition. To reduce this noise, we try to remove tags that are only weakly connected to the other tags in their community. In principle, SBMs can identify such nodes using MCMC simulations. However, the size of our question-tag matrix renders this approach computationally infeasible. Therefore, we here use a simpler approach.
In particular, we construct a weighted unipartite tag network, where edge weights denote the intensity with which tags co-occur on the same questions. Next, we label nodes in this tag network according to the community structure we discovered above. For each tag, we now calculate the overrepresentation, $O_{tc}$, of (weighted) links to nodes in each of the 247 communities:
@@
$O_{tc}$ measures how important tag $t$ is to community $c$ compared to all other communities. High values of $O_{ct}$ indicate that the tag is a core part of the community, while low $O_{ct}$ values signal that the tag may be misplaced or only weakly related to community $c$. To focus on core tags, for each community, we remove the $20\%$ of tags with the lowest $O_{ct}$ values. Finally, we also remove communities with fewer than 3 tags, for which the tag composition provides too little information on the associated skill. In this way, we arrive at a skill space with 237 skills, composed of a total of 4,054 tags.
Appendix B Skill space
In this section, we provide a detailed, annotated version of the skill space depicted in Fig. 1b. This annotated skill space is shown in Fig. B.1.
Each node is labeled with an example of a programming skill provided by ChatGPT-4.0. To do so, we prompt the LLM as follows:
I have a list of tags related to software engineering and computer science. The tags are ordered by their importance, listing the most important tags first. Using no more than 10 words, could you please suggest a general task related to the following tags ’task_description’ as key words? Please make sure that the task is general and not too specific. Please only return json content and not anything else.
Because questions describe hands-on coding problems, this prompt refers to tasks (generalized problems), which we regard as closely related to the skills they require. Doing so avoids that the LLM tries to guess which skills are required to solve the problem stated in a question.
Figure B.1: Fully annotated skill space.
Fig. B.2a illustrates how SO activity shifted across software expertise areas between 2009 and 2022. Notable growth appears in Web Frameworks, certain Android-related skills, Cloud Computing, and AI/ML. Conversely, skills associated with Advanced Programming Concepts (e.g., skills to Develop a data structure library for efficient data manipulation, Develop a multimedia application in Windows, or Develop a thread-safe application to prevent race conditions), as well as to Networking or Enterprise and Web App Development, have seen their relative importance on SO decline.
Figure B.2: Skill dynamics. a. Skill user-share change from 2009 to 2022. Orange markers signal increases, purple markers decreases, in user-shares between 2009 and 2022. Marker transparency reflects the size of shifts: darker tones indicate larger changes in user shares.